OpenStudy (anonymous):

Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). A. drawing a black two out of a standard 52 card deck given it’s not a face card or an Ace. B. drawing a six out of a standard 52 card deck given it’s not a face card or an Ace. C. rolling a sum of 7 on two fair six sided dice D. rolling a sum of 3 on two fair six sided dice E. rolling a sum of 10 on two fair six sided dice

OpenStudy (anonymous):

this is maybe the tenth time i have seen this strange question where does it come from? you have 5 probabilities to compute, so it is a pain

OpenStudy (anonymous):

ughhh it's something crazy I just can't seem to understand it

OpenStudy (anonymous):

A has probability \(\frac{1}{2}\) since half the cards are black B has probability \(\frac{4}{36}=\frac{1}{9}\) since you have 6 sixes in a deck without the 16 face cards or aces

OpenStudy (anonymous):

I got A correct but when it comes to the other ones I don't understand it very well

OpenStudy (anonymous):

there are six ways to get a 7 on two dice, 36 possible rolls, so probability of C is \(\frac{6}{36}\)

OpenStudy (anonymous):

which ones are not clear?

OpenStudy (anonymous):

Algebra: Probability and statistics Solvers Lessons Answers archive Quiz In Depth If you need immediate math help from PAID TUTORS right now, click here. (paid link) Click here to see ALL problems on Probability-and-statistics Question 386490: Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, seperated by a comma. A standard deck of cards has 12 face cards and four aces (aces are not face cards). Please show your probability for each choice below. a) drawing a red five out of a standard 52 card deck given it's not a face card or an ace. b) rolling a sum of 10 on two fair six sided dice. c) rolling a sum of 3 on two fair six sided dice. d) rolling a sum of 7 on two fair six sided dice. e) drawing a four out of a standard 52 card deck given it's not a face card or an ace. (Scroll Down for Answer!) Did you know that Algebra.Com has hundreds of free volunteer tutors who help people with math homework? Anyone can ask a math question, and most questions get answers! OR get immediate PAID help on: Go!!! Answer by ewatrrr(10589) (Show Source): You can put this solution on YOUR website! Hi, a) drawing a red five out of a standard 52 card deck given it's not a face card or an ace. 52 - 4aces - 12face cards = 36 possibilities P = 2/36 = 1/18 e) drawing a four out of a standard 52 card deck given it's not a face card or an ace. 4/36 = 1/9 b) rolling a sum of 10 on two fair six sided dice. P = 3/36 = 1/12 c) rolling a sum of 3 on two fair six sided dice. P = 2/36 = 1/18 d) rolling a sum of 7 on two fair six sided dice. P = 6/36 = 1/6 a,c have equal probabilities of success ___1_2_3__4__5__6 1|_2_3_4__5__6__7 2|_3_4_5__6__7__8 3|_4_5_6__7__8__9 4|_5_6_7__8__9_10 5|_6_7_8__9_10_11 6|_7_8_9_10_11_12

OpenStudy (anonymous):

wow, that is very helpful, ty so much

OpenStudy (anonymous):

is just the whole concept of the problem that I get a little confuse

OpenStudy (anonymous):

the number of ways to roll a total of ten on two dice \((4,6),(6,4),(5,5)\) there are 3 ways out of 36 probability is \(\frac{3}{36}\) similarly there are 2 ways to roll a total of 3 \((1,2),(2,1)\) so that probability is \(\frac{2}{36}\)

OpenStudy (anonymous):

where the hell did that come from?

OpenStudy (anonymous):

letter C. 6/36=1/6??? do I have it right?

OpenStudy (anonymous):

lol IDK

OpenStudy (anonymous):

can you see what I c

OpenStudy (anonymous):

aah i see it is a black two for A there are two of them, out of the 36 remaining cards so that probability is \(\frac{2}{36}\) same as the probability of rolling a 3 on two dice

OpenStudy (anonymous):

Hahahaha that was some GREAT TROLLING.

OpenStudy (anonymous):

the question asks for the two that are the same your answer is A and D

OpenStudy (anonymous):

your answer is A and D

OpenStudy (anonymous):

lol yes it was

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

question asks "List the letters of the two choices below which have equal probabilities "

OpenStudy (anonymous):

uh? are you giving me a ? if you are where are the choices

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

ty all

OpenStudy (anonymous):

i think c or d